OpenStudy (anonymous):

Write the equation of a circle that has a center at the origin and a radius with a length of 4 inches

OpenStudy (anonymous):

@M_lowreen help me plz

OpenStudy (bradely):

Standard equation of circle with center (0,0) is x^2 +y^2 =r^2 Here r is the radius Substitute r value to get equation Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

so it x2+y2=4

OpenStudy (bradely):

yes